OpenStudy (anonymous):

Look at this expression. PLEASE IM SO CONFUSED :( x^5y^3z^3 over xy^-3z^-1 (im supposed to find the simplest form of this expression)

jimthompson5910 (jim_thompson5910):

when dividing terms like this, you subtract the corresponding exponents

jimthompson5910 (jim_thompson5910):

x^5 divided by x = x^5 divided by x^1 = x^(5 - 1) = x^4 So x^5 divided by x = x^4 Do the same for the other terms

OpenStudy (campbell_st):

the problem appears to be that the denominator involves negative powers... \[\frac{x^5y^3z^3}{xy^{-3}z^{-1}}\] just use the rule shown above but take care when subtracting a negative power.... which occurs with y and z

jimthompson5910 (jim_thompson5910):